OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π) 2 cos x + sin 2x = 0

OpenStudy (anonymous):

hi kcb,, the trig identity is: sin 2x=2sin x cos x

OpenStudy (anonymous):

therefore, 2cos x +2sin x cos x+0

OpenStudy (anonymous):

im sorry,its equal to zero 2cos x +2sin x cos x=0

OpenStudy (anonymous):

now take out the cos x,... cos x (2+2sin x)=0

OpenStudy (anonymous):

here both are equaled to zero that is, cos x =0 and 2+2 sin x = 0

OpenStudy (anonymous):

x=arc cos 0 and 2sin x=-2 sin x=-2/2 sin x=-1 x=arc sin -1 x=pi/2 and x=-pi/2